40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Feb 28 20:49:33 2012, kiwamu, Update, LSC, insalling OSA 
    Reply  Wed Feb 29 04:23:14 2012, kiwamu, Update, LSC, REFL OSA installed 
       Reply  Mon Mar 5 05:39:36 2012, kiwamu, Update, LSC, REFL OSA installed APtable.png
          Reply  Mon Mar 5 13:11:06 2012, kiwamu, Update, LSC, installed a long BNC cable for REFL OSA 
Message ID: 6352     Entry time: Mon Mar 5 05:39:36 2012     In reply to: 6340     Reply to this: 6354
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: REFL OSA installed 

The OSA for the REFL beam is now fully functional.

The only thing we need is a long BNC cable going from the AP table to the control room so that we can monitor the OSA signal with an oscilloscope.

The attached picture shows how they look like on the AP table. Both AS and REFL OSAs are sitting on the corner region.

Quote from #6340

I placed the OSA (Optical Spectrum Analyzer) on the AP table and this OSA will monitor the REFL beam.

Tomorrow I will do fine alignment of the OSA.

Attachment 1: APtable.png  2.672 MB  | Hide | Hide all
ELOG V3.1.3-